OpenStudy (trojanpoem):

If I is the electric current, A the area of the conductor, the number of electrons for the unit of mass is n_o the speed of this electrons is v so the equation I = 0.25 n_oeva is

OpenStudy (trojanpoem):

@Michele_Laino

OpenStudy (michele_laino):

the relòationship between the current density \(j\) and the number of electrons, is: \(j=nev\), where \(e=1.6 \cdot 10^{-19} coulombs\)

OpenStudy (michele_laino):

relationship*

OpenStudy (michele_laino):

\(n \) is the number of electrons over volume

OpenStudy (michele_laino):

and \(j=I/A\)

OpenStudy (michele_laino):

I don't understand your coefficient 0.25

OpenStudy (trojanpoem):

It's written like this in the question

OpenStudy (trojanpoem):

so the unit of n_o is 1/kg in (SI) ?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

from my steps, i can write this: \(I=n_oevA\)

OpenStudy (trojanpoem):

M.L^2T^-2* M^-1* L * T^-2 * L^2 L^4 T^-4 ?

OpenStudy (trojanpoem):

I can't be right in anyway.

OpenStudy (michele_laino):

I'm thinking at your formula, and it can not be correct if \(a\) is the area and \(e\) is the electron charge since we have this: right side= \[\frac{{coulombs}}{{Kg}} \cdot \frac{m}{{\sec }} \cdot {m^2}\]

OpenStudy (michele_laino):

which is not a current

OpenStudy (trojanpoem):

coulombs ?

OpenStudy (michele_laino):

yes! I measure \(e\) in coulombs

OpenStudy (trojanpoem):

I meant, why don't you replace it with A * S

OpenStudy (trojanpoem):

A * T

OpenStudy (michele_laino):

A is the cross sectional area of your conductor

OpenStudy (trojanpoem):

Yeah.

OpenStudy (michele_laino):

\(v\) is the average speed of conduction electrons, so we have: \(I=0.25n_0\;e\;v\;A=\) \[0.25 \cdot \frac{{coulombs}}{{Kg}} \cdot \frac{m}{{\sec }} \cdot {m^2}\]

OpenStudy (michele_laino):

is 0.25 a quantity with a specific unit of measure?

OpenStudy (michele_laino):

for example, can 0.25 be a density?

OpenStudy (trojanpoem):

In dimension problems, It's not necessary to bring to us an equation which we know. Today was the first time to see this. but I think 0.25 is just a constant like last 2 pi

OpenStudy (trojanpoem):

Do you suspect that 0.25 has a unit ?

OpenStudy (michele_laino):

yes! since the right side of your formula, is not a current, whereas if 0.25 is a density, then its unit of measure is \(Kg/m^3\), so we can write this: \[\begin{gathered} 0.25 \cdot \frac{{coulombs}}{{Kg}} \cdot \frac{m}{{\sec }} \cdot {m^2} = \frac{{Kg}}{{{m^3}}} \cdot \frac{{coulombs}}{{Kg}} \cdot \frac{m}{{\sec }} \cdot {m^2} = \hfill \\ \hfill \\ = \frac{{coulombs}}{{\sec }} = amperes \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

furthermore if \(\rho=m/V\), then we can write this: \[\frac{{dn}}{{dm}} = \frac{{dn}}{{\rho dV}} = \frac{1}{\rho }\frac{{dn}}{{dV}}\]

OpenStudy (trojanpoem):

What does dn/dm mean?

OpenStudy (michele_laino):

it is the number of electrons over mass of conductor

OpenStudy (trojanpoem):

I looks like a derivative xD

OpenStudy (michele_laino):

yes! whereas \(dn/dV\) is my density of electrons, namely number of electrons over volume of conductor

OpenStudy (trojanpoem):

You said: "since the right side of your formula, is not a current, whereas if 0.25 is a density, then its unit of measure is Kg/m3" Well that's convincing but how to ensure so ? the question is as written here.

OpenStudy (michele_laino):

It was my hypothesis, since it is the first time that I see your formula

OpenStudy (michele_laino):

when I was at university, I have always used this one: \[j = {n_0}ev\]

OpenStudy (trojanpoem):

choices: a) correct,b) correct according to dimensions c) wrong d)wrong " " "

OpenStudy (michele_laino):

I think it is wrong!

OpenStudy (trojanpoem):

So it's wrong :) thanks again again

OpenStudy (michele_laino):

:)

OpenStudy (michele_laino):

I understand the presence of 0.25

OpenStudy (michele_laino):

if we have this formula: \[I = {n_0}evA\] with \(n_0\) as number of electrons over mass, we can say that it is wrong according to the dimensions. nevertheless, due to the presence of the coefficient 0.25, we have to say that your formula is wrong, since using dimensional analysis we are not able to write numerical coefficients

OpenStudy (michele_laino):

here is why there is the coefficient 0.25

OpenStudy (trojanpoem):

So 0.25 makes the formula wrong not wrong according to the dimensions ?

OpenStudy (michele_laino):

yes!

OpenStudy (trojanpoem):

I tried googling this formula and no results

OpenStudy (michele_laino):

please wait

OpenStudy (michele_laino):

I'm sorry, as I wrote before, using units of measure, the right side is not a current, so I think, that your formula is wrong according to dimensions

OpenStudy (trojanpoem):

so when to say wrong, when to say wrong according to dimensions

OpenStudy (michele_laino):

I say wrong according dimensions when, as in our case, the right has the wrong dimensions

OpenStudy (michele_laino):

right side*

OpenStudy (trojanpoem):

so wrong when the constant is wrong ?

OpenStudy (michele_laino):

yes I think so!

OpenStudy (trojanpoem):

But there is no way to know that In the last question you said correct as you knew that T= 2 pi sqrt(L/g) you know the formula but I don't.

OpenStudy (michele_laino):

for example if I write this formula: \[T = 3\sqrt {\frac{l}{g}} \] what can we conclude?

OpenStudy (trojanpoem):

Wrong, but if It's the first time to see it ?

OpenStudy (trojanpoem):

wait wait, in Problem #21 , Prove that I = n_oeva according to dimensions

OpenStudy (trojanpoem):

See the 6th formula in this table: https://quizlet.com/3348670/physics-formulae-and-constants-flash-cards/

OpenStudy (michele_laino):

it is wrong according to dimensions, since N is the number of free electrons/m^3, whereas A is the cross sectional area so m^2

OpenStudy (trojanpoem):

I made a mistake :o

OpenStudy (trojanpoem):

I have combined two parts of two questions

OpenStudy (michele_laino):

no worries! :)

OpenStudy (trojanpoem):

n is the number of electrons for the unit of volume , SORRY

OpenStudy (michele_laino):

and A is the cross sectional area whose unit of measure is m^2

OpenStudy (trojanpoem):

so it's I = neva L^-3 * A * T * L * T^-1 * L^2 = A

OpenStudy (trojanpoem):

so as you said it's wrong

OpenStudy (trojanpoem):

What's the difference between G , g ?